A $98,000 trust is to be invested in bonds paying 6%, CDs paying 5%, and mortgages paying 7%. The sum of the bond and CD investment must equal the mortgage investment. to earn $6130 annual income from the investment, how much should the bamk invest in bonds?
Answer by venugopalramana(3286)

A $98,000 trust is to be invested in bonds paying 6%, CDs paying 5%, and mortgages paying 7%. The sum of the bond and CD investment must equal the mortgage investment. to earn $6130 annual income from the investment, how much should the bamk invest in bonds?
LET BONDS INVESTMENT=B
...CD.....=C
....MORTGAGE...=M
B+C+M=98000........I
B+C=M........B+C-M=0......II
EQN.I-EQN.II
2M=98000
M=49000
B+C=49000.......B=49000-C.................III
INTEREST ON BONDS =B*6/100=0.06B
....CD...=0.05C
.....M.....=0.07M
TOTAL INTEREST =0.06B+0.05C+0.07M=6130.........IV
0.06(49000-C)+0.05C+0.07*49000=6130
0.01C=49000(0.06+0.07)-6130=6370-6130=240
C=24000
B=49000-24000=25000